# Rapid Rotation of Polarization Orientations in PSR B1919$+$21’s Single Pulses: Implications On Pulsar’s Magnetospheric Dynamics

### v1

@@ Abstract
We analyse and model rapid rotations of polarization orientations in PSR B1919+21’s single pulses, based on FAST observation data. In more than one third of B1919+21’s single pulses, polarization angle (PA) is found to rotate quasi-monotonically with pulse longitude, by over 180 or even 360 degrees. Most of these quasi-monotonic PA curves have negative slopes with respect to pulse longitude. Oscillations of circular polarization fraction accompany these PA rotations. This rapid rotation could be induced by quick change of phase lag between normal wave modes within an individual pulse. We propose a phenomenological model to successfully reproduce the observed polarization rotations in single pulses, and calculate phase lags in a dipolar magnetic field of an aligned rotating pulsar, with a dispersion relation of orthogonal wave modes in strongly magnetized plasma. The weak frequency dependence of observed polarization rotation requires the radio emission height to be low, approximately 10 times the neutron star radius.
@@ 1 Introduction
The highly magnetized relativistic plasma spheres around neutron stars, i.e. magnetospheres, generate and modify pulsars’ radio emission (see Philippov & Kramer, 2022 for a review). Magnetospheres not only play a vital role in pulsar and neutron star theories, but are also natural laboratories for particle experiments (e.g., Noordhuis et al., 2023; Xue et al., 2023).
@@
Among theoretical, especially analytical studies, much effort has been focused on analyzing normal wave modes in magnetospheres. A fundamental property of the magnetosphere is the birefringence, so dispersion relations of the ordinary (O) mode and the extraordinary (E/X) mode are the basis for further analysis of wave propagation. Melrose & Stoneham (1977) derived the dispersion relations of O and E modes in a cold relativistic plasma with a single bulk velocity value and infinitely large magnetic field strength, and naturally explained pulsars’ orthogonal polarization modes observed by Manchester et al. (1975). Such dispersion relations have been applied to describe ray trajectories in pulsar magnetospheres (e.g., Barnard & Arons, 1986; Petrova & Lyubarskii, 2000) and to explain the origin of circular polarization in pulsar radio emissions (e.g., Lyubarskii & Petrova, 1998), involving calculating the phase lag between wave modes.
@@
Despite the success of models, some parameters in pulsar magnetospheres are not well-constrained. Two key parameters are the plasma multiplicity $\kappa$, defined as the ratio between the actual plasma density and the Goldreich-Julian particle number density (equation (9) in Goldreich & Julian, 1969), and the Lorentz factor of the “secondary” plasma particles, $\gamma_{\mathrm{sec}}$. The “secondary” plasma particles are electron positron pairs produced in polar cap pair cascade processes. The Lorentz factor of these particles is also related to the emission height, the distance between emission point and pulsar centroid. Usually $\kappa$ takes the value $10^{3}\sim 10^{5}$, and $\gamma_{\mathrm{sec}}$ takes the value $10\sim 10^{3}$ (e.g., Philippov & Kramer, 2022). Beside some attempts to achieve more realistic simulations, studying observations’ connection with theories is important for pushing studies forward. The double pulsar system J0737--3039 has provided good observational constraints on the magnetospheric physics (Lyutikov & Thompson, 2005; Lower et al., 2024). However, for solitary pulsars, detailed studies of individual pulse behaviours might bring us more insight into pulsar magnetospheres, with China’s the Five-hundred-meter Aperture Spherical radio Telescope (FAST) now providing a large amount of high-quality radio pulsar data.
@@
In this letter we present our analysis and modeling of some rapid polarization position angle (PA) rotations in PSR B1919$+$21’s single pulses, based on FAST data. As the first pulsar ever discovered (Hewish et al., 1968), B1919$+$21 is frequently studied under various frequency bands. It is a typical normal pulsar with $P=1.34$ s and $\dot{P}=1.35\times 10^{-15}$ (Manchester et al., 2005). B1919$+$21 exhibits sub-pulse drifting, with drifting properties changing with frequency (Cordes, 1975; Proszynski & Wolszczan, 1986). It was also found to exhibit unusual radius-frequency-mapping (RFM), with its pulse profile width slightly getting wider when frequency increases (Mitra & Rankin, 2002; Hassall et al., 2012; Pilia et al., 2016). Besides, polarization studies on this pulsar have shown that the single pulses’ PA distribution changes significantly with frequency (Mitra et al., 2015; Olszanski et al., 2019). Dyks (2017, 2019) modelled the 45 degree PA jumps observed in B1919+21 (Mitra et al., 2015) through the coherent superposition of orthogonally polarized proper mode waves. Primak et al. (2022) have identified a slow rotation of polarization direction within a limited longitude range. The rotations occurred on the time scale of subpulse drift and led to a torus-shaped distribution of polarization states in Stokes space. Our work will further investigate single pulses’ polarization patterns, and try to relate them to theoretical issues in pulsar magnetospheres.
@@
In Section 2, basic information on observation and data reduction is given. Section 3 shows the observed results ranging from basic properties of the integrated profile and the highly linearly polarized profile, to single pulses’ distribution patterns, and finally the rapid rotating polarization orientations in individual single pulses. In Section 4 we develop the model to explain the quick rotating single pulse polarization orientations in two steps: firstly a phenomenological model for reproducing the observed polarization patterns and exploring the parameter space, secondly the normal modes’ phase lag calculation in a more realistic magnetosphere, with known dispersion relation. We propose that the key factor to understand the polarization pattern is the phase lag value between orthogonal modes. Discussions on model feasibility, B1919$+$21’s precursor component, asymmetries in B1919$+$21’s polarization pattern and possible particularity of B1919$+$21 are presented in Section 5. Section 6 summarizes the conclusions.
@@ 3.1 Overview of integrated profiles’ properties
The integrated profile of all pulses is shown in Figure 1, where the total intensity ($I$), linear polarization intensity ($L$), circular polarization intensity ($V$), PA ($\psi$) and ellipticity angle (EA, $\chi$) at each longitude are plotted. There is a precursor component on the leading edge of profile, which contributes to the long extension of PA and EA curves to the left. Definitions of PA and EA are in Eq. A12 and A13. The polarization profiles are in accord with the results in Weisberg et al. (1999). B1919$+$21’s polarization is complex, with a PA curve that vastly deviates from Rotating Vector Model (RVM) curves (Radhakrishnan & Cooke, 1969). Except for longitudes around the leading and trailing edges of the pulse profile, the linear polarization fraction ($L/I$) is small ($<0.2$).
@@
Figure 1: (i) Integrated profile of 2947 pulses of B1919$+$21. Black line - total intensity ($I$); Cyan line - 10 times total intensity ($10\times I$); red dashed line - linear polarization intensity ($L=\sqrt{Q^{2}+U^{2}}$); blue line - circular polarization intensity ($V$). $I$, $L$ and $V$ are normalized by the maximum total intensity of the respective profiles. Green dots in $\psi$ panel with errorbars - polarization position angles (PA, $\psi=0.5\arctan(U/Q)$). Magenta dots in $\chi$ panel with errorbars - ellipticity angel (EA, $\chi=0.5\arcsin(V/{\sqrt{Q^{2}+U^{2}+V^{2}}})$). (ii) Intensity profiles at different observation frequencies. 38 MHz: Pilia et al. (2016); 119 MHz: Bilous et al. (2016); 326 MHz: Olszanski et al. (2019); 610 MHz: Gould & Lyne (1998); 1250 MHz: this paper; 1642 MHz: Gould & Lyne (1998); 4850 MHz: Hoensbroech (1999). All higher-frequency profiles are shifted in longitude to be roughly aligned (achieve minimum residual between profiles) with the 38 MHz profile. (iii) A zoomed in version of the left plot with a $L/I$ panel. Three vertical grey dashed lines represent the longitudes $-0.925^{\circ}$, $-0.457^{\circ}$ and $0.875^{\circ}$. (iv) A zoomed in version of the middle plot. In the $\psi$ panel, a purple dotted curve and a yellow dotted curve are added, which represents the PA curve shifted by $45^{\circ}$ and $90^{\circ}$. Only longitudes where $L/\sigma_{L}>10$ are chosen for plotting ($\sigma_{L}$ is the standard deviation of $L$ in off-pulse longitude range).
@@
In the longitude range of the leading component, the PA curve shows three “discontinuities”, which are marked with three vertical grey dashed lines in Figure 1. The first one is a nearly $90^{\circ}$ jump at longitude $-0.925^{\circ}$, accompanied with a local minimum in $L$ and $L/I$, and a sign change of $V$ and $\chi$. The second one is a $45^{\circ}$ ($135^{\circ}$, most clear in Figure 1 (iii)) jump at longitude $-0.457^{\circ}$ with a local minimum in $L$ and $L/I$, and a local maximum in $|V|$ and $|\chi|$. The third one is a nearly $45^{\circ}$ mild jump at longitude $0.875^{\circ}$ with a local minimum in $L$ and $L/I$, and a local maximum in $|V|$ and $|\chi|$. After the third “discontinuity”, the PA curve is continuous but distorted, characteristic with low $L/I$, in the central region of the pulse profile. The $V$ and $\chi$ curves are also complex. The first “discontinuity” could be interpreted as the change of dominant mode between two orthogonal polarization modes (OPMs), whose existence has been reported by former observations (e.g., Mitra et al., 2015). The third one has also been recorded and discussed before (e.g., Mitra et al., 2015; Dyks, 2017), which is associated with coherent summation of OPMs. The second one, however, seems newly discovered, and may have an origin similar to the third “discontinuity”.
We’ve also added our intensity profile into a sequence of pulse profiles under different frequencies, shown in Figure 1 (iv). This plot confirms the profiles’ widening when frequency increases. The references for other frequencies’ profiles are provided in the caption.
@@ 3.2 Attempts at deriving radiative geometry
To derive radiative geometry, we adopt the method that is put forward in Mitra et al. (2023) and tested in Johnston et al. (2024). Practically, at each longitude, we choose all samples having linear polarization fraction $L/I>0.8$ and PA measurement error $\sigma_{\mathrm{PA}}<5^{\circ}$, to add them up and form a highly linearly polarized profile. The result is shown in Figure 2. Comparing with the total integrated profile in Figure 1, highly linearly polarized signals only appear at on-pulse window’s edges, and represent only one of the OPMs. We tried RVM curve fitting of the PA dots in the highly linearly polarized profile: for each pair of ($\alpha$,$\zeta$), we calculate the least chi-square value when changing ($\psi_{0}$, $\phi_{0}$) ($\phi_{0}$ is limited between $-90^{\circ}$ and $90^{\circ}$). The fitting process is same as that in Cao et al. (2024). The minimum chi-squares plot is also presented in Figure 2. For B1919$+$21, the inclination angle $\alpha$ is poorly constrained, but the impact angle $\beta=\zeta-\alpha$ seems constrained to a small value, which is consistent with the estimation by Kuz’min & Wu (1992) based on the radiation cone model. In all plots, the $\phi=0^{\circ}$ longitude is chosen as the best fitting of $\phi_{0}$, which is near the leading component peak’s longitude.
Figure 2: Left: profile from adding only highly linearly polarized pulse bins (for details please refer to the text in Section 3.2). The lowest panel: the polarization profile, lines’ meanings are same as Figure 1. The other 3 panels: distributions of $L/I$, PA ($\psi$) and EA ($\chi$) versus longitude. The blue dots with errorbars in PA and EA panels are directly calculated (through Eq. A12 and A13) from the highly polarized pulse profile in the lowest panel. Right: the common logarithm values of chi-square (divided by degree of freedom) of fitting RVM curve of given ($\alpha$,$\zeta$) to the modified PA curve. The red “+” sign marks the ($\alpha$,$\zeta$) where we get the smallest $\chi^{2}$.
@@ 3.3 Single pulses’ polarization distributions
Firstly we calculate the 2D histogram of $L/I$, PA and EA for all pulses in the on-pulse window, and the result is shown in Figure 3. In the PA panel, we directly see enhancements of single-pulse PA distributions that represent two OPMs, around the longitude range of the first two vertical grey dashed lines. On the right side of the third dashed line, even the individual pulses’ PA distribution doesn’t seem to follow RVM-like tracks, though OPMs are likely to also appear at longitude $3\sim 4^{\circ}$. The EA’s distribution forms several tracks, varying from positive values to negative values.
Figure 3: (i) Distributions of PA, EA and $L/I$ versus longitude, of all 2 observation epochs’ 2947 pulses. Panel 4 (1-4 from up to bottom): the integrated profiles of all pulses taken into account, where lines’ meanings are same as Figure 1; Panel 3: the distributions of $L/I$; Panel 2: the distributions of PA. Panel 1: the distributions of EA. For panel 1, 2 and 3, the distribution is more concentrated when color turns darker. In all panels, three vertial grey lines are the same as that of Figure 1. (ii) Distribution of circular polarization fraction painted on $\phi-\psi$ plot (for calculation details please refer to the text). (iii) Similar plot for intensity distribution (normalized with the maximum intensity value in the distribution). Only single pulses’ bins where both $\sigma_{\text{PA}}\leq 5^{\circ}$ and $\sigma_{\text{EA}}\leq 5^{\circ}$ are included, to make sure that all presented points are with significant linear and circular polarization. The blue dots with errorbars are calculated from the integrated profile, same as Figure 1.
The histograms in Figure 3 (i) are made by counting the number of pulses within the longitude and PA bins. For a given longitude $\phi$ and a given PA $\psi$, we calculate the average circular polarization fraction ($\bar{V}/\bar{I}$) and average intenity $\bar{I}$ of samples whose longitudes are within ($\phi$, $\phi+\Delta\phi$) and PAs are within ($\psi$, $\psi+\Delta\psi$). We choose $\Delta\phi=360^{\circ}/4096$, and $\Delta\psi=1^{\circ}$. The histograms of the average circular polarization fraction ($\bar{V}/\bar{I}$) and average intenity $\bar{I}$ are shown in Figure 3 (ii) (iii) respectively. The $V/I$ distribution plot shows the difference in circular polarization senses for OPMs around $\phi=-1^{\circ}$, which is natural for normal wave modes in magnetized plasmas (e.g., Beskin & Philippov, 2012; Cao et al., 2024). Moreover, in the central region ($1^{\circ}<\phi<8^{\circ}$), the $V/I$ sense distribution is related to the integrated profile’s PA curve: in longitude range ($1^{\circ}$, $3^{\circ}$), the integrated profile’s PAs correspond to $V>0$; while in longitude range ($6^{\circ}$, $8^{\circ}$), the integrated PA curve seems to be located around the $V=0$ region in the histogram for individual pulse bins. Such distributions should have some physical implications.

### v2

@@ Abstract
We analyze and model rapid rotations of polarization orientations in PSR B1919$+$21’s single pulses based on Five-hundred-meter Aperture Spherical radio Telescope observation data. In more than one-third of B1919$+$21’s single pulses, the polarization position angle (PA) is found to rotate quasi-monotonically with pulse longitude, by over 180 or even 360. Some single pulse PA even rotates by over 540. Most of these quasi-monotonic PA curves have negative slopes with respect to pulse longitude. Oscillations of circular polarization fraction accompany these PA rotations. This rapid rotation could be induced by a quick change of phase lag between two normal wave modes within an individual pulse. We propose a phenomenological model to reproduce the observed polarization rotations in single pulses, and calculate phase lags in a dipolar magnetic field of an aligned rotating pulsar, with a dispersion relation of orthogonal wave modes in strongly magnetized electron-positron plasma. According to the dispersion relation, the weak frequency dependence of observed polarization rotation requires small angles between the radio wavevector and local magnetic fields, which requires the radio emission height to be low, on the order of 10 times neutron star radius.
@@ 1 Introduction
Magnetospheres around neutron stars generate and modify pulsar radio emission (see Philippov & Kramer, 2022 for a review). Magnetospheres not only play a vital role in pulsar and neutron star theories, but are also natural laboratories for particle physics (e.g., Noordhuis et al., 2023; Xue et al., 2023).
@@
Among theoretical studies, especially analytical studies, much effort has been focused on analyzing normal wave modes in magnetospheres. A fundamental property of the magnetosphere is the birefringence, so dispersion relations of the ordinary (O) mode and the extraordinary (X/E) mode are the basis for further analysis of wave propagation. Melrose & Stoneham (1977) derived the dispersion relations of O and X modes in a cold relativistic plasma with a single bulk velocity value and infinitely large magnetic field strength, and naturally explained pulsars’ orthogonal polarization modes observed by Manchester et al. (1975). Such dispersion relations have been applied to describe ray trajectories in pulsar magnetospheres (e.g., Barnard & Arons, 1986; Petrova & Lyubarskii, 2000) and to explain the origin of circular polarization in pulsar radio emissions (e.g., Lyubarskii & Petrova, 1998), involving calculating the phase lag between wave modes.
@@
Despite the success of models, some parameters in pulsar magnetospheres are not well constrained. Two key parameters are the plasma multiplicity $\kappa$, defined as the ratio between the actual plasma density and the Goldreich-Julian particle number density (Equation (9) in Goldreich & Julian, 1969), and the Lorentz factor of the secondary plasma particles, $\gamma_{\mathrm{sec}}$. The secondary plasma particles are electron positron pairs produced in polar cap pair cascade processes. The Lorentz factor of these particles is also related to the emission height, the distance between the emission point, and the pulsar centroid. Usually $\kappa$ takes the value $10^{3}\sim 10^{5}$, and $\gamma_{\mathrm{sec}}$ takes the value $10\sim 10^{3}$ (e.g., Philippov & Kramer, 2022). Beside some attempts to achieve more realistic simulations, studying observations’ connection with theories is important for pushing studies forward. The double pulsar system J0737--3039 has provided good observational constraints on the magnetospheric physics (Lyutikov & Thompson, 2005; Lower et al., 2024a). However, for solitary pulsars, detailed studies of individual pulse behaviors might bring us more insight into pulsar magnetospheres, with China’s Five-hundred-meter Aperture Spherical radio Telescope (FAST) now providing a large amount of high-quality radio pulsar data.
@@
In this paper, we present our analysis and modeling of some rapid polarization position angle (PA) rotations in PSR B1919$+$21’s single pulses, based on FAST data. As the first pulsar ever discovered (Hewish et al., 1968), B1919$+$21 is frequently studied under various frequency bands. It is a typical normal pulsar with spin period $P=1.34$ s and period derivative $\dot{P}=1.35\times 10^{-15}$ (Manchester et al., 2005). B1919$+$21 exhibits sub-pulse drifting, with drifting properties changing with frequency (Cordes, 1975; Proszynski & Wolszczan, 1986). It was also found to exhibit unusual radius-frequency mapping (RFM), with its pulse profile width slightly increasing as the frequency increases (Mitra & Rankin, 2002; Hassall et al., 2012; Pilia et al., 2016). Besides, polarization studies on this pulsar have shown that the PA distribution of single pulses changes significantly with frequency (Mitra et al., 2015; Olszanski et al., 2019). Dyks (2017, 2019) modeled the 45 PA jumps observed in B1919+21 (Mitra et al., 2015) through the coherent superposition of orthogonally polarized proper-mode waves. Primak et al. (2022) have identified a slow rotation of polarization direction within a limited longitude range. The rotations occurred on the time scale of subpulse drift and led to a torus-shaped distribution of polarization states in Stokes parameter space ($Q/I,U/I,V/I$). Our work will further investigate single pulses’ polarization patterns, and try to relate them to theoretical issues in pulsar magnetospheres.
@@
In Section 2, basic information on observation and data reduction is given. Section 3 shows the observed results, ranging from the basic properties of the integrated profile and the highly linearly polarized profile, to single pulses’ distribution patterns, and finally, the rapid rotating polarization orientations in individual single pulses. In Section 4 we develop the model to explain the quick rotating single-pulse polarization orientations in two steps: firstly a phenomenological model for reproducing the observed polarization patterns and exploring the parameter space, secondly the normal modes’ phase lag calculation in a more realistic magnetosphere, with known dispersion relation. We propose that the key factor for understanding the polarization pattern is the phase lag value between orthogonal modes. Discussions on model feasibility, possible alternative models, and asymmetries in B1919$+$21’s polarization pattern are presented in Section 5. Section 6 summarizes the conclusions.
@@ 2 Observation and data reduction
$L=\left\{\begin{aligned} &\sqrt{Q^{2}+U^{2}-\epsilon_{L}},&Q^{2}+U^{2}-\epsilon_{L}>0\\ &0,&Q^{2}+U^{2}-\epsilon_{L}<0\end{aligned}\right.$ (1)
where
$\epsilon_{L}=\dfrac{Q^{2}\sigma_{U}^{2}+U^{2}\sigma_{Q}^{2}}{Q^{2}+U^{2}}$ (2)
Similarly, for calculating total polarization intensity $P$, we have
$P=\left\{\begin{aligned} &\sqrt{Q^{2}+U^{2}+V^{2}-\epsilon_{P}},&Q^{2}+U^{2}+V^{2}-\epsilon_{P}>0\\ &0,&Q^{2}+U^{2}+V^{2}-\epsilon_{P}<0\end{aligned}\right.$ (3)
where
$\epsilon_{P}=\dfrac{Q^{2}(\sigma_{U}^{2}+\sigma_{V}^{2})+U^{2}(\sigma_{V}^{2}+\sigma_{Q}^{2})+V^{2}(\sigma_{Q}^{2}+\sigma_{U}^{2})}{Q^{2}+U^{2}+V^{2}}$ (4)
Errors of $L$ and $P$ could be derived from equations above with error propagation. We skip the details here.
@@ 3.1 Overview of integrated profiles’ properties
The integrated profile of all pulses is shown in Figure 1, where the total intensity ($I$), linear polarization intensity ($L$), circular polarization intensity ($V$), PA ($\psi$), and ellipticity angle (EA, $\chi$) at each longitude are plotted. There is a precursor component on the leading edge of the profile, which contributes to the long extension of PA and EA curves to the left. Definitions of PA and EA are in Equation A12. The error of PA could then be derived from error propagation:
$\sigma_{\mathrm{PA}}=\dfrac{1}{2}\dfrac{\sqrt{Q^{2}\sigma_{U}^{2}+U^{2}\sigma_{Q}^{2}}}{Q^{2}+U^{2}}$ (5)
The EA from the data is derived as
$\mathrm{EA}=0.5\arctan(V/L)$ (6)
where $L$ is defined in Equation 1. And the error of EA is similar to that of PA:
$\sigma_{\mathrm{EA}}=\dfrac{1}{2}\dfrac{\sqrt{L^{2}\sigma_{V}^{2}+V^{2}\sigma_{L}^{2}}}{L^{2}+V^{2}}$ (7)
The polarization profiles are in accord with the results in Weisberg et al. (1999). B1919$+$21’s polarization is complex, with a PA curve that vastly deviates from rotating vector model (RVM) curves (Radhakrishnan & Cooke, 1969). Except for longitudes around the leading and trailing edges of the pulse profile, the linear polarization fraction ($L/I$) is small ($<0.2$).
@@
Figure 1: (i) Integrated profile of 2947 pulses of B1919$+$21. Black line—total intensity ($I$); Cyan line—10 times total intensity ($10\times I$); red dashed line—linear polarization intensity ($L=\sqrt{Q^{2}+U^{2}}$); blue line—circular polarization intensity ($V$). $I$, $L$ and $V$ are normalized by the maximum total intensity of the respective profiles. Red dots in $L/I$ panel with errorbars—linear polarization degrees. Green dots in $\psi$ panel with errorbars—polarization position angles (PA, $\psi=0.5\arctan(U/Q)$). Magenta dots in $\chi$ panel with errorbars—ellipticity angel (EA, $\chi=0.5\arcsin(V/{\sqrt{Q^{2}+U^{2}+V^{2}}})$). A zoomed in version of the $I$, $L$, and $V$ profiles are added as the second panel (from bottom to top). Three vertical gray dashed lines represent the longitudes $-0^{\circ}_{.}925$, $-0^{\circ}_{.}457$ and $8750^{\circ}_{.}$ 875. (ii) A more zoomed in version of $I$, $L$, $V$ profiles, $L/I$, $\psi$ and $\chi$ panels. In the $\psi$ panel, a purple dotted curve and a yellow dotted curve are added, which represents the PA curve shifted by $45^{\circ}$ and $90^{\circ}$. Only longitudes where $L/\sigma_{L}>10$ are chosen for plotting ($\sigma_{L}$ is the standard deviation of $L$ calculated from Equation 1 through error propagation).
@@
In the longitude range of the leading component, the PA curve shows three “discontinuities”, which are marked with three vertical gray dashed lines in Figure 1. The first one is a nearly $90^{\circ}$ jump at longitude $-0^{\circ}_{.}925$, accompanied by a local minimum in $L$ and $L/I$, and a sign change of $V$ and $\chi$. The second one is a $45^{\circ}$ ($135^{\circ}$, most clear in Figure 1 (ii)) jump at longitude $-0^{\circ}_{.}457$ with a local minimum in $L$ and $L/I$, and a local maximum in $|V|$ and $|\chi|$. The third one is a nearly $45^{\circ}$ mild jump at longitude $8750^{\circ}_{.}$ 875 with a local minimum in $L$ and $L/I$, and a local maximum in $|V|$ and $|\chi|$. After the third “discontinuity”, the PA curve is continuous but distorted, characteristic with low $L/I$, in the central region of the pulse profile. The $V$ and $\chi$ curves are also complex. The first “discontinuity” could be interpreted as the transition between two orthogonal polarization modes (OPMs), whose existence has been reported by previous observations (e.g., Mitra et al., 2015). The third one has also been recorded and discussed before (e.g., Mitra et al., 2015; Dyks, 2017), which is associated with coherent summation of OPMs. The second one, however, seems newly discovered, and may have an origin similar to the third “discontinuity”.
@@ 3.2 Attempts at deriving radiative geometry
To derive radiative geometry, we adopt the method proposed in Mitra et al. (2023) and tested in Johnston et al. (2024). Practically, at each longitude, we choose all samples having linear polarization fraction $L/I>0.8$ and PA measurement error $\sigma_{\mathrm{PA}}<5^{\circ}$, to add them up and form a highly linearly polarized profile. The result is shown in Figure 2. Compared with the total integrated profile in Figure 1, highly linearly polarized signals appear only at the edges of the on-pulse window, and represent only one of the OPMs. The highly linearly polarized profile has two emission components at the leading edge of the total integrated profile. The two components are close to each other, with a large depth in between. This phenomenon is similar to the bifurcated emission component studied in Dyks et al. (2010); Dyks (2023). We tried RVM curve fitting of the PA dots in the highly linearly polarized profile: for each pair of ($\alpha$,$\zeta$), we calculate the least chi-square value when changing ($\psi_{0}$, $\phi_{0}$) ($\phi_{0}$ is limited between $-90^{\circ}$ and $90^{\circ}$). The fitting process is the same as that in Cao et al. (2024). The minimum chi-squares plot is also presented in Figure 2. For B1919$+$21, the inclination angle $\alpha$ is poorly constrained, but the impact angle $\beta=\zeta-\alpha$ seems constrained to a small value, which is consistent with the estimate by Kuz’min & Wu (1992) based on the radiation cone model. In all plots, the $\phi=0^{\circ}$ longitude is chosen as the best-fitted $\phi_{0}$, which is near the longitude of the leading component peak.
Figure 2: Left: profile from adding only highly linearly polarized pulse bins (for details, please refer to the text in Section 3.2). $L$ bias is also corrected here. The lowest panel: the polarization profile, the meanings of the lines are the same as those in Figure 1. The other three panels: distributions of $L/I$, PA ($\psi$) and EA ($\chi$) versus longitude. The blue dots with error bars in the PA and EA panels are directly calculated (through Equations A12 and A13) from the highly polarized pulse profile in the lowest panel. Right: the common logarithm values of chi-square (divided by degree of freedom) of fitting the RVM curve of given ($\alpha$,$\zeta$) to the modified PA curve. The red “+” sign marks the ($\alpha$,$\zeta$) where we get the smallest $\chi^{2}$.
@@ 3.3 Single pulses’ polarization distributions
First, we calculate the 2D histogram of $L/I$, PA, and EA for all pulses in the on-pulse window, and the result is shown in Figure 3 (i). The histograms in Figure 3 (i) are made by counting the number of pulses within the longitude and PA bins. For a given longitude $\phi$ and a given PA $\psi$, we calculate the average circular polarization fraction ($\bar{V}/\bar{I}$) and the average intensity $\bar{I}$ of samples whose longitudes are within ($\phi$, $\phi+\Delta\phi$) and PAs are within ($\psi$, $\psi+\Delta\psi$). We choose $\Delta\phi=360^{\circ}/4096$, and $\Delta\psi=1^{\circ}$. The histograms of the average circular polarization fraction ($\bar{V}/\bar{I}$) and the average intensity $\bar{I}$ are shown in Figure 3 (ii) and (iii) respectively. In the PA panel, we directly see enhancements of single-pulse PA distributions that represent two OPMs, around the longitude range of the first two vertical gray dashed lines. On the right side of the third dashed line, even the individual pulses’ PA distribution does not seem to follow RVM-like tracks, though OPMs are likely to also appear at longitude $3^{\circ}\sim 4^{\circ}$. The EA’s distribution forms several tracks, varying from positive values to negative values.
Figure 3: (i) Distributions of PA, EA, and $L/I$ versus longitude of all two observation epochs’ 2947 pulses. Panel 4 (1-4 from top to bottom): the integrated profiles of all pulses taken into account, where the meanings of the lines are the same as Figure 1; panel 3: the distributions of $L/I$; panel 2: the distributions of PA; panel 1: the distributions of EA. For panel 1, 2 and 3, the distribution is more concentrated when the color turns darker. In all panels, three vertial gray lines are the same as those in Figure 1. (ii) Distribution of circular polarization fraction painted on $\phi-\psi$ plot (for calculation details, please refer to the text). (iii) Similar plot for intensity distribution (normalized with the maximum intensity value in the distribution). Only single pulses’ bins where both $\sigma_{\text{PA}}\leq 5^{\circ}$ and $\sigma_{\text{EA}}\leq 5^{\circ}$ are included to make sure that all presented points are with significant linear and circular polarization. The blue dots with error bars are calculated from the integrated profile, the same as in Figure 1.
The $V/I$ distribution plot shows the difference in circular polarization senses for OPMs around $\phi=-1^{\circ}$, which is natural for normal wave modes in magnetized plasmas (e.g., Beskin & Philippov, 2012; Cao et al., 2024). Moreover, in the central region ($1^{\circ}<\phi<8^{\circ}$), the $V/I$ sense distribution is related to the integrated profile’s PA curve: in longitude range ($1^{\circ}$, $3^{\circ}$), the integrated profile’s PAs correspond to $V>0$; while in longitude range ($6^{\circ}$, $8^{\circ}$), the integrated PA curve seems to be located around the $V=0$ region in the histogram for individual pulse bins. Such distributions should have some physical implications.